# thanks: Current affiliation: ISIS Neutron and Muon Source, STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, UK

### v1

@@ VI Conclusion
In this work, we have performed Monte-Carlo ray-tracing simulations of several multiplexed triple-axis spectrometer models relying on prismatic analysis to obtain a quasi-continuous energy-transfer coverage with only one or two incident energies $E_{i}$. Building on the concept and the design of the CAMEA spectrometer at PSI, we have introduced a distinct data analysis approach relying on the statistical likelihood that a neutron detected on a given pixel corresponds to a particular final energy $E_{f}$. This approach is dubbed Positionally-Calibrated Prismatic Analysis (PCPA). Simulations of the spectrometer model for a simple inelastic scattering signal show that PCPA data analysis maximizes the number of final energies measured simultaneously. Departing from the CAMEA design with eight analyzer stations, we explored further design evolutions, including changing the number of analyzer stations and the crystal mosaic of the graphite analyzers. These conceptual studies highlight the enormous flexibility offered by prismatic analysis. In principle, any existing TAS with a PSD located far enough from the analyzer to allow for distance collimation effects to occur can make use of the prismatic analyzer concept and the PCPA technique. Our conceptual work serves as a foundation for performing more comprehensive studies of multiplex prismatic spectrometers, taking collimation and background mitigation into account. This includes testing a more complex sample kernel, considering thermal diffuse scattering from the analyzers, and dealing with complex sample environments. Our work utilized a realistic primary spectrometer developed in Ref. [19], including a fully optimized neutron guide system for the cold source of the High Flux Isotope Reactor. Although no formal concept has been chosen for the secondary spectrometer, when combined with a refined, CAMEA-style multiplexed prismatic system, a concept known as MANTA, our work shows that this modern beam-line would deliver world-class performance to study excitations in quantum condensed matter physics.
@@ Acknowledgements.
The work at Georgia Tech was supported by the Department of Energy, Basic Energy Sciences, Neutron Scattering Program under grant DE-SC-0018660. Initial work by A.D. was supported by the Georgia Institute of Technology’s Letson Fellowship and the President’s Undergraduate Research Award. We are grateful for the fruitful discussions with Mark Lumsden, Lowell Crow Jr., Jaime Fernandez-Baca, Barry L. Winn, Michael Hoffmann, Ian Turnbull, Mads Bertleson, Daniel Mazzone, and Jonas Birk that helped advance this work. We are indebted to Felix Groitl for providing is with the McStas model of CAMEA.

### v2

@@ VI Conclusion
In this work, we have performed Monte-Carlo ray-tracing simulations of several multiplexed triple-axis spectrometer models relying on prismatic analysis to obtain a quasi-continuous energy-transfer coverage with only one or two incident energies $E_{i}$. Building on the concept and the design of the CAMEA spectrometer at PSI, we have introduced a distinct data analysis approach relying on the statistical likelihood that a neutron detected on a given pixel corresponds to a particular final energy $E_{f}$. This approach is dubbed Positionally-Calibrated Prismatic Analysis (PCPA). Simulations of the spectrometer model for a simple inelastic scattering signal show that PCPA data analysis maximizes the number of final energies measured simultaneously. Departing from the CAMEA design with eight analyzer stations, we explored further design evolutions, including changing the number of analyzer stations and the crystal mosaic of the graphite analyzers. These conceptual studies highlight the enormous flexibility offered by prismatic analysis. In principle, any existing TAS with a PSD located far enough from the analyzer to allow for distance collimation effects to occur can make use of the prismatic analyzer concept and the PCPA technique. Our conceptual work serves as a foundation for performing more comprehensive studies of multiplexed prismatic spectrometers, taking collimation and background mitigation into account. This includes testing a more complex sample kernel, considering thermal diffuse scattering from the analyzers, and dealing with complex sample environments. Our work utilized a realistic primary spectrometer developed in Ref. [19], including a fully optimized neutron guide system for the cold source of the High Flux Isotope Reactor. Although no formal concept has been chosen for the secondary spectrometer, when combined with a refined, CAMEA-style multiplexed prismatic system, a concept known as MANTA, our work shows that this modern beam-line would deliver world-class performance to study excitations in quantum condensed matter physics.
@@ Acknowledgements.
The work at Georgia Tech was supported by the Department of Energy, Basic Energy Sciences, Neutron Scattering Program under grant DE-SC-0018660. Initial work by Adit S. Desai was supported by the Georgia Institute of Technology’s Letson Fellowship and the President’s Undergraduate Research Award. The work of Marcus Daum was supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education for the DOE under contract number DE‐SC001466. This work used resources of the High Flux Isotope Reactor at Oak Ridge National Laboratory which is a DOE office of Science User Facility. We are grateful for the fruitful discussions with Mark Lumsden, Lowell Crow Jr., Jaime Fernandez-Baca, Barry L. Winn, Michael Hoffmann, Ian Turnbull, Mads Bertleson, Daniel Mazzone, and Jonas Birk that helped advance this work. We are indebted to Felix Groitl for providing us with the McStas model of CAMEA.